### v4

# Potential-Barrier Affinity Effect in Solid Systems

Abstract
Electron accumulation in interatomic regions is a fundamental quantum phenomenon dictating chemical bonding and material properties, yet its origin remains elusive across disciplines[1, 2, 3, 4]. Here, we report a quantum accumulation effect—potential-barrier affinity (PBA)—revealed by solving the Schrödinger equation for a crystalline potential. PBA effect drives significant interatomic electron accumulation when electron energy exceeds the barrier maximum. This effect essentially enhances interatomic electron density, governing microstructures and properties of condensed matter. Our theory overturns the traditional wisdom that the interstitial electron localization in electride requires potential-well constraints[2, 3, 4] or hybrid orbitals[5], and it serves as the fundamental mechanism underlying the formation of conventional solid bonding. This work delivers a paradigm shift in understanding electron distribution and establishes a theoretical foundation for the microscopic design of material properties.
Introduction
As one of the most fundamental quantum phenomena, the complex behavior of electrons in interatomic region enables diverse bonding types, which govern the structures and properties of materials, making it a cornerstone of condensed matter physics, theoretical chemistry, and materials science. In particular, the interatomic electron accumulation with diverse density patterns—such as the conventional covalent bonds[6] and the unique interstitial anion electrons (IAEs) found in electrides[7, 8, 9, 3, 4]—provides a versatile platform for regulating and designing material properties. For instances, diamond, the hardest known natural material, is renowned for its strong covalent bonding feature[10]; the first room-temperature air-stable inorganic electride, Ca24Al28O64[11], was experimentally confirmed to exhibit the IAEs, with a measured superconducting transition temperature of 0.4 K[12]; and electrides Cs3O and Ba3N possess one-dimensional IAEs together with strong interrod interactions induce band inversion in a two-dimensional superatomic triangular lattice, resulting in Dirac node lines[13]. Although these interatomic electron accumulation patterns govern remarkable functional properties in materials, the fundamental origin of chemical bonding remains a subject of active debate[14, 15, 16, 17, 18]. Emerging research suggests it constitutes a complex electronic behavior governed by a confluence of mechanisms, including node-induced electron confinement, Pauli repulsion, orbital contraction, and polarization[1]. In electrides, conventional understanding holds two divergent views on the source of IAEs: the first explains them as confined bound states within interatomic potential wells[2, 3, 4, 9], and the second as a consequence of multicentered bonding induced by the overlap of hybrid orbitals[5].
Results
Abnormal interstitial electron accumulation
Figure 1: Structure and electronic localization function of Na-hP4. a, Structure of Na-hP4 (space group P63/mmc) at 320 GPa. Lattice parameters are $a=b=2.784$ Å and $c=3.873$ Å with two inequivalent atomic positions of $2a(0,0,0)$ and $2d(2/3,1/3,1/4)$. b, Electron localization function of Na-hP4 plotted in the (110) plane at 320 GPa.
Figure 2: Evidence of the existence of near-free-valence electrons. a, The Kohn-Sham effective local potential in the (110) plane of Na-hP4. b, Band structure (left) and density of states (right) of Na-hP4. c, The effective local potential $V_{eff}^{KS}$, total electron density($\rho$), bound-state derived electron density ($\rho_{B}$), and near-free-electron density ($\rho_{F}$) along $\bar{1}1$] direction of Na-hP4, where “grey coverage” denotes the interatomic region. d, Energy differences between Fermi energy level and maximum point of Kohn-Sham effective local potential ($E_{F}-V_{M}$) for seven typical electrides.
The transparent dense sodium has been experimentally confirmed and is recognized as a typical electride[19, 4], which has attracted much attention due to the anomalous metal-insulator phase transition under high pressure. In this study, we conducted first-principles calculations on the electronic structure of Na-hP4 at 320 GPa[19], as shown in Figure 1a, to reveal its anomalous characteristics, including the presence of a finite bandgap ($\sim$1.75 eV) and the formation mechanisms of IAEs. The electron localization function depicted in Figure 1b clearly shows the highly localized features of the electrons in the interstitial regions, which are consistent with the results of the previously reported differential charge density[19]. The computational details can be found in Methods section.
We subsequently calculate the effective local potential by the Kohn-Sham density functional theory[20] in Figure 2a, along with the valence band maximum (VBM) and conduction band minimum (CBM). Contrary to the conventional expectation of potential wells, we found that electrons localize at the maximum point of the potential ($V_{M}$) within the interstitial region. Furthermore, we observed that the Fermi energy level ($E_{F}=16.0$ eV) is much higher than the maximum point of the potential ($V_{M}=4.4$ eV), indicating that near-free electrons (NFEs) occupy the single-particle states with energy levels ranging from $V_{M}$ to $E_{F}$. The band structure shown in Figure 2b (left) displays two valence electron bands (4 electrons per cell) located within the NFE energy range. This suggests that all valence electrons behave as unbound-state NFEs rather than being confined as interstitial bound electrons. While early models interpreted the interstitial-potential well as arising from Pauli exclusion by core electrons[21, 2], a quantitatively accurate description of the Pauli potential remains absent within the fermion quasiparticle framework. We address this point further in the Method section.
Additionally, the density of states (DOS) in Figure 2b (right) demonstrates that the s, p, and d components are negligible compared to the total DOS within the valence-electron energy range, which is also a distinct feature commonly associated with unbound-state NFEs[22], rather than the multicentered bonding of hybrid orbitals. In addition, the concept of hybrid orbital derives from a linear combination of localized atomic orbitals centered on a single atom site[23]. As such, it should be applied with caution when describing unbound states. To further clearly demonstrate the presence of the electron localization within the interstitial region, we plot the Kohn-Sham effective local potential ($V_{eff}^{KS}$), total ($\rho$), bound-state ($\rho_{B}$), and near-free ($\rho_{F}$) electron densities along the $\bar{1}1$] direction of Na-hP4 in Figure 2c. We observed that the NFE bands contribute almost all the interstitial electron density near the maximum potential point, while the electron density tails from the bound-state bands can be ignored. This result directly confirms that interstitial electrons originate from near-free-valence states. We also present the effective ionic potential of Na+-hP4 in Figure 5 of the Method Section by removing the valence electrons. The analysis further confirms that IAEs behave as NFEs with high density accumulation around the potential maximum. Notably, we further found evidence for the existence of NFEs ($E_{F}>V_{M}$) in representative electrides[24, 25, 8, 26, 27, 28, 29], as illustrated in Figure 2d, suggesting that the existence of near-free valence electrons is a common phenomenon among electrides. More evidence of the NFE-density distribution in the interstitial region can be found in Figure 6 and Figure 7 of the Method section. These findings compel a reevaluation of the current understanding of IAEs in electrides.
Figure 3: Unbound-state Kronig-Penney model. a, Periodic square potential well and barrier model, where “A” and “B” denote the atomic and interatomic regions; $V_{0}$” is the height of the square potential barrier; “aa” and “bb” are the potential-well and primary-cell width, respectively. b, near-free-electron band structure of Kronig-Penney model. c, Absolute wavefunctions of $n=1$ and $n=2$ bands. d, Electron density calculated for $n=1$ band ($\rho_{1}$) and $n=2$ band ($\rho_{2}$), as well as the total density $\rho=\rho_{1}+\rho_{2}$, where “grey coverage” denotes the “B” region. The used parameters of the Kronig-Penney model are a=1.0a=1.0Å, b=6.185b=6.185Å, and $V_{0}=114.0$ eV.
Potential-barrier affinity effect
To explore the nature of NFE states ($\epsilon_{nk}>V_{0}$) in a periodic field, as shown in Figure 3a, we establish a one-dimensional generalized Kronig-Penney (KP) model for unbound states, providing a direct counterpart to the traditional bound-state scheme[30], where $\epsilon_{nk}$ indicates the eigenvalue of $nk$th Bloch state ($\psi_{nk}$); “A” and “B” denote the atomic and interatomic regions, respectively. The detailed information of KP model in this work can be found in Method Section. According to the Schrödinger equation, the wavefunctions of unbound states have a general form:
$\psi_{nk}^{\text{A}}(x)=A_{nk}e^{i\alpha_{nk}x}+A^{\prime}_{nk}e^{-i\alpha_{nk}x},\penalty 10000\ x\in\text{A};$
$\psi_{nk}^{\text{B}}(x)=B_{nk}e^{i\beta_{nk}x}+B^{\prime}_{nk}e^{-i\beta_{nk}x},\penalty 10000\ x\in\text{B},$ (1)
where $\psi^{\text{A}}_{nk}$ and $\psi^{\text{B}}_{nk}$ are the branches of the wavefunction in regions A and B, respectively; $A_{nk}$, $A_{nk}^{\prime}$, $B_{nk}$, and $B_{nk}^{\prime}$ are the corresponding plane-wave-component coefficients in the wavefunction. The wavevectors ($\alpha_{nk}$ and $\beta_{nk}$) in A and B regions are $\alpha_{nk}=\sqrt{2\epsilon_{nk}}$ and $\beta_{nk}=\sqrt{2(\epsilon_{nk}-V_{0})}$, respectively. The details of the numerical solution for the Schrödinger equation of KP model can be found in the Method Section. In Figure 3b, we present the unbound-state band structure ($\epsilon_{nk}>V_{0}$) under the KP potential, which shows finite band gaps and parabolic characteristics as the perturbed free-electron gas[31, 32]. The concavities of the unbound-state bands in Figure 3b are also similar to those in Figure 2b for Na-hP4 near the $G$ point.
Subsequently, we calculated the absolute value of the wavefunctions and the electron density in Figure 3c and Figure 3d, respectively, corresponding to the $n=1,2$ bands in Figure 3b. All of them exhibit counterintuitive phenomena, where the large amplitudes are observed in the interatomic region. This is consistent with the electron density of Na-hP4 in Figure 2c. Furthermore, we reproduced the double peaks of the interatomic electron density in Figure 2c by our unbound-state KP model. Figure 3c shows that the band index ($n$) is directly related to the number of peaks of $|\psi_{nk}(x)|$ in the barrier region. Herein, we do not consider the electron density in the atomic region because of the large differences in the atomic central potential and the square potential. Instead, we will mainly use the unbound-state KP model to demonstrate that the steep-shaped potential can lead to a high electron density in the interatomic region. In the unbound-state regime ($\epsilon_{nk}>V_{0}$), we can directly find that the wavefunction exhibits the properties of short and long wavelengths in regions A (atomic) and B (interatomic) regions, respectively, because of the inequality $\alpha_{nk}>\beta_{nk}$. Therefore, the wavefunction-continuity conditions at the point of barrier-well junction, such as $x=0$, can be written as:
$\psi_{nk}^{\text{A}}(0)=\psi_{nk}^{\text{B}}(0);$
$\left.\frac{d\psi_{nk}^{\text{A}}}{dx}\right|_{x=0}=\left.\frac{d\psi_{nk}^{\text{B}}}{dx}\right|_{x=0}.$ (2)
The continuity conditions in Eq. (2) ensure the wavefunction smoothness at $x=0$. In particular, the first-derivative continuity of the wavefunction leads to the amplitude of the long-wavelength plane-wave component being greater than that of the short-wavelength one, resulting in the PBA effect with high electron density distribution in the interatomic region. Quantitatively, the following equations can be derived for the coefficients by Eqs. (12):
$B_{nk}=\frac{A_{nk}+A_{nk}^{\prime}}{2}+\frac{1}{\sqrt{1-V_{0}/\epsilon_{nk}}}\frac{A_{nk}-A_{nk}^{\prime}}{2}=\bar{A}_{nk}+\lambda_{nk}\Delta_{nk};$
$B_{nk}^{\prime}=\frac{A_{nk}+A_{nk}^{\prime}}{2}-\frac{1}{\sqrt{1-V_{0}/\epsilon_{nk}}}\frac{A_{nk}-A_{nk}^{\prime}}{2}=\bar{A}_{nk}-\lambda_{nk}\Delta_{nk}.$ (3)
With Eq. (3), we found that the coefficient $B$ and $B^{\prime}$ are determined by the average $\bar{A}_{nk}=(A_{nk}+A_{nk}^{\prime})/2$ and difference $\Delta_{nk}=(A_{nk}-A_{nk}^{\prime})/2$ of coefficients $A_{nk}$ and $A_{nk}^{\prime}$, as well as the factor $\lambda_{nk}=(\sqrt{1-V_{0}/\epsilon_{nk}})^{-1}$. For the special case, $\Delta_{nk}=0$ for $A_{nk}=A_{nk}^{\prime}$, the plane-wave components share the same amplitudes $|A_{nk}|=|A_{nk}^{\prime}|=|B_{nk}|=|B_{nk}^{\prime}|$. For the general cases, $\Delta_{nk}\neq 0$, $V_{0}/\epsilon_{nk}$ is an important quantity that determines the amplitudes of the plane-wave components. When the energy level is much higher than the potential barrier ($\epsilon_{nk}>>V_{0}$, or $\lambda_{nk}\to 1$), the difference in the component amplitudes between A and B regions can be ignored; Conversely, if the energy level approaches the height of the potential barrier ($\epsilon_{nk}\to V_{0}$ and $\lambda\to\infty$), the unbound-sate wavefunction exhibits a high component characteristic in the barrier region:
$\lim_{\epsilon_{nk}\to V_{0}}|B_{nk}|,|B_{nk}^{\prime}|\sim\lambda_{nk}|\Delta_{nk}|>>|A_{nk}|,|A_{nk}^{\prime}|.$ (4)
So far, we demonstrate that the wavefunctions derived from Eq. (4), with high-component amplitudes within interatomic region, will result in a high interatomic electron density (PBA effect). Therefore, we immediately draw the conclusion: if there are sufficiently large occupied states (or DOS) near the potential barrier ($V_{0}$), the system can exhibit excellent PBA electrons, such as the IAEs in electrides. Here, $V_{0}$ can be regarded as $V_{M}$ in Figure 2a when the lowest potential point is shifted to 0 eV.
According to the results of the KP model, we find that the PBA of quantum state is dramatically consistent with the classical physical intuition: a free particle has an initial kinetic energy $T_{0}$; When it passes through the barrier region $V(x)>0$, part of the kinetic energy is converted into potential energy, and its velocity decreases due to the smaller kinetic energy $T(x)=T_{0}-V(x)<T_{0}$, causing it to have a higher probability of being detected in the higher barrier region. When the initial kinetic energy is much higher than the potential barrier $T_{0}>>V(x)$, the particle can move freely throughout space $T(x)\approx T_{0}$. This finding seems to indicate that high-energy PBA electrons tend to classical behavior according to Bohr’s correspondence principle[33].
[…]
Metallic and covalent bonds
[…]
Discussion
In summary, we theoretically reveal a universal PBA effect by solving the Schrödinger equation of a one-dimensional crystal potential. The PBA effect is characterized by significant electron accumulation within the interatomic region, arising from the smooth connection of wavefunctions between short-wavelength components in the atomic region and long-wavelength components across the interatomic region, amplifying the latter and producing electron accumulation there. We demonstrate that the PBA effect becomes most pronounced as the electron energy approaches the barrier maximum and fundamentally governs the resulting electron density distribution in condensed matter, thereby dictating their microstructures and properties. From the PBA effect, we derive the necessary conditions for the existence of electrides: (i) the Fermi level lies above the maximum of the effective potential barrier; (ii) the barrier maximum is located in a spatially open region far from nuclei; and (iii) an appreciable density of occupied states exists between the Fermi level and the barrier maximum. Such conditions provide a physically transparent guideline for the discovery of an electride. Furthermore, this theory overturns the traditional belief that the formation of IAEs strictly requires potential well constraints or multicentered bonding, and it serves as the fundamental mechanism underlying the formation of conventional solid bonding, defect states, and excited-state electron density distribution. This unexpected discovery represents a fundamental paradigm shift in understanding electron behavior in materials, providing a universal theoretical foundation for the microscopic design of specific material properties.
Methods
First-principles calculations
The first-principles electronic structures, including effective local potential, electron localization function, electron density, and band structure, are carried out using the Kohn-Sham density functional theory (KSDFT)[20] as implemented in the Vienna Ab initio Simulation Package (VASP v6.4.1)[34]. The Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional[35] and the projected-augmented-wave method[36] are adopted for all KSDFT calculations. The kinetic energy cutoff of 1000 eV for the plane-wave basis set and $k$-spacing of 0.15 Å-1 within Monkhorst-Pack sampling[37] are used for all cases. The dense real-space grids of $120\times 120\times 160$ are adopted for calculating the high-resolution electron localization function in Figure 1b. Notably, the crystal structures of Na[19], Ba2N[24], Ca2N[8], Li5Si[26], Li6Al[27], Li8Au[28] , Li4Rh[29] utilized in this study were obtained from previously published sources; the structure of diamond and Al come from the built-in structure library of Material Studio.
Figure 5: The Kohn-Sham effective potential and electron density of Na+-hP4 in (a) (110) plane, and (b) $\bar{1}1$] direction. The grey line denotes the reference potential of Na-hP4.
To isolate the interatomic potential field from the screening effect of valence electrons, we calculated the KSDFT-derived effective potential for a hypothetical system of Na+-hP4 by removing the valence electrons out of Na-hP4. As shown in Figure 5(a) and (b), the maximum of this “ion-only” potential remains located at the interstitial site. This is expected, as the potential converges toward the superposition of screened ionic Coulomb potentials in the interatomic region. Notably, the absence of near-free electron screening leads to a higher interatomic potential barrier in Na+-hP4 than in Na-hP4. Crucially, this independent calculation confirms our central thesis: interstitial anionic electrons (IAEs) are near-free electrons with high density accumulation around the potential maximum, not bound states confined to potential wells.
Near-free electrons in electrides
Figure 6: The electron density distribution functions for near-free electrons ($\rho_{F}$) and bound-state electrons ($\rho_{B}$), where $\rho_{F}$ and $\rho_{B}$ are calculated according to the unbound-state bands ($V_{M}<\epsilon_{nk}<E_{F}$) and bound-state bands ($\epsilon_{nk}<V_{M}$), respectively. The isosurfaces of 0.008, 0.005, 0.008, 0.01, 0.04, 0.1, and 0.05 Å-3 are used for Ca2N, Ba2N, Sr2N, Li5Si, Li6Al, Li8Au, and Li4Rh, respectively.
Figure 7: (a) The electron density distribution functions for near-free electrons ($\rho_{F}$), and (b) the effective potential and electron density calculated along the dashed line of the Ca-Ca pair for Ca24Al28O64, where the isosurface is 0.001 Å-3.
To further establish that IAEs arise from near-free-electron (NFE) behavior, we calculated the partial electron density from band decomposition for the electrides in Fig. 2(d) of the maintext, as well as the representative mayenite [Ca24Al28O64][11]. The results, presented in the Figure 6 and Figure 7, reveal that the NFE density accumulates predominantly in the interstitial region, in agreement with the findings for Na-hP4.
Unbound-state solver for Kronig-Penney potential
Solving the unbound states within the KP potential is the foundation for understanding the anomalous PBA effect. Here, we provide the complete details to solve the Schrödinger equation:
$\frac{1}{2}\frac{d^{2}\psi_{nk}(x)}{dx^{2}}+[\epsilon_{nk}-V(x)]\psi_{nk}(x)=0,$ (5)
where $V(x)$ is the periodic KP potential[30] as shown in Figure 3a. The unbound-state wavefunction can be written in a direct sum of regions A and B:
$\psi_{nk}(x)=\psi_{nk}^{\text{A}}(x)\oplus\psi_{nk}^{\text{B}}(x),$ (6)
where the formal solutions of $\psi_{nk}^{\text{A}}$ and $\psi_{nk}^{\text{B}}$ can be found in Eq. (1). We then can obtain the wavefunction with the periodicity of the primitive cell according to Bloch’s theorem[38] ($u_{nk}=e^{-ikx}\psi_{nk}$):
$u_{nk}^{\text{A}}(x)=e^{-ikx}\psi_{nk}^{\text{A}}(x)=A_{nk}e^{i(\alpha_{nk}-k)x}+A_{nk}^{\prime}e^{-i(\alpha_{nk}+k)x},\penalty 10000\ x\in\text{A};$
$u_{nk}^{\text{B}}(x)=e^{-ikx}\psi_{nk}^{\text{B}}(x)=B_{nk}e^{i(\beta_{nk}-k)x}+B_{nk}^{\prime}e^{-i(\beta_{nk}+k)x},\penalty 10000\ x\in\text{B},$ (7)
where the wavevectors $\alpha_{nk}=\sqrt{2\epsilon_{nk}}$ and $\beta_{nk}=\sqrt{2(\epsilon_{nk}-V_{0})}$ represent the short and long wavelengths, respectively, as mentioned in the Results section. To obtain the coefficients ($A_{nk}$, $A_{nk}^{\prime}$, $B_{nk}$, $B_{nk}^{\prime}$) and eigenvalues ($\epsilon_{nk}$), we applied the continuity conditions at $x=0$ and the periodic boundary conditions at $x=a$ and x=(ba)x=-(b-a):
$u_{nk}^{\text{A}}(0)=u_{nk}^{\text{B}}(0),$
$\left.\frac{du_{nk}^{\text{A}}}{dx}\right|_{x=0}=\left.\frac{du_{nk}^{\text{B}}}{dx}\right|_{x=0};$
$u_{nk}^{\text{A}}(a)=u_{nk}^{\text{B}}(-(b-a)),$
$\left.\frac{du_{nk}^{\text{A}}}{dx}\right|_{x=a}=\left.\frac{du_{nk}^{\text{B}}}{dx}\right|_{x=-(b-a)}.$ (8)
By combining Eqs. (7), and (8), we can derive the following linear equations:
$\left(\begin{array}[]{cccc}1&1&-1&-1\\ \alpha_{nk}-k&-(\alpha_{nk}+k)&-(\beta_{nk}-k)&\beta_{nk}+k\\ e^{i(\alpha_{nk}-k)a}&e^{-i(\alpha_{nk}+k)a}&-e^{-i(\beta_{nk}-k)(b-a)}&-e^{i(\beta_{nk}+k)(b-a)}\\ (\alpha_{nk}-k)e^{i(\alpha_{nk}-k)a}&-(\alpha_{nk}+k)e^{-i(\alpha_{nk}+k)a}&-(\beta_{nk}-k)e^{-i(\beta_{nk}-k)(b-a)}&(\beta_{nk}+k)e^{i(\beta_{nk}+k)(b-a)}\end{array}\right)\left(\begin{array}[]{c}A_{nk}\\ A_{nk}^{\prime}\\ B_{nk}\\ B_{nk}^{\prime}\end{array}\right)=0.$ (9)
If Eq. (9) has the nonzero vector solutions, the determinant of the $\times4$ matrix must be zero, which leads to the following equation:
$cos(kb)=-\frac{\alpha^{2}+\beta^{2}}{2\alpha\beta}sin(\alpha a)sin(\beta(b-a))+cos(\alpha a)cos(\beta(b-a)).$ (10)
[…]
In this work, the parameters of the square potential barrier in KP model are adopted as close as possible to the first-principles results of Na-hP4: $V_{0}=114.0$ eV and $a=1.0$ Å are obtained from the height and half width, respectively, of $V_{eff}^{KS}$ near the atomic region; $b=6.185$ Å is set as the length of $\bar{1}1$] direction of Na-hP4 according to Figure 2c. 21 $k$ points are used for band structure and electron density calculations in the KP model.
Generalized doulbe-barrier Kronig-Penney model
Figure 8: (a) The effective potential and electron density calculated by the all-electron potential for Diamond; (b) The generalized KP model with two different barriers (or wells) in a unit cell. “A”/“C” and “B”/“D” denote the interatomic regions and atomic regions, respectively. Especially, “C” also indicates the covalent bonding region.
Figure 4(c) of the maintext shows that the effective potential in diamond reaches a maximum at the carbon sites. This feature originates from the pseudopotential representing the core-electron interaction, an inherent characteristic of the frozen-core approximation within the PAW method. To clarify the physical picture, we performed all-electron (AE) calculations, the results of which are presented in Figure 8(a). These calculations confirm the highly localized core-electron density at the carbon nuclei and the corresponding deep effective potential wells centered on these atomic sites.
To give a physical picture of the covalent bonding [see Figure 4(c) in our maintext], we further propose a more generalizable extension of the Kronig-Penney (KP) model, which can be implemented as a double-barrier (or double-well) in a unit cell for the diamond structure, as illustrated in Figure 8(b). Here, we shift the maximum potential to 0 eV for convenience.
For such a square barrier-well potential, the components of the wavefunction adopt one of two different forms depending on the electron energy: (i) an oscillatory, planewave-like solution when the energy exceeds the potential barrier, or (ii) an exponentially decaying solution when the energy is below the potential barrier. For the square barrier-well potential with the configuration of $V_{B}\leq V_{D}<V_{C}<V_{A}$ as shown in Figure 8(b), the energy levels ($\epsilon_{nk}$) can be classified into three categories: (1) bound states ($V_{B}<\epsilon_{nk}<V_{C}$) for localized electrons (core states); (2) semibound states ($V_{C}<\epsilon_{nk}<V_{A}$) for covalent bonds; and (3) unbound states ($\epsilon_{nk}>V_{A}$) for near-free electrons, metallic bonds, or interstitial anionic electrons.
For comparison with the covalent bonds in the diamond system, our analysis focuses on the semibound state in Figure 8(b) with energy in between the two barriers ($V_{C}<\epsilon_{nk}<V_{A}$ ), where the wavefunction components in different regions can be written as:
$\psi_{nk}^{\text{A}}(x)=A_{nk}e^{\alpha_{nk}x}+A^{\prime}_{nk}e^{-\alpha_{nk}x},\penalty 10000\ x\in\text{A};$
$\psi_{nk}^{\text{B}}(x)=B_{nk}e^{i\beta_{nk}x}+B^{\prime}_{nk}e^{-i\beta_{nk}x},\penalty 10000\ x\in\text{B};$
$\psi_{nk}^{\text{C}}(x)=C_{nk}e^{i\gamma_{nk}x}+C^{\prime}_{nk}e^{-i\gamma_{nk}x},\penalty 10000\ x\in\text{C};$
$\psi_{nk}^{\text{D}}(x)=D_{nk}e^{i\eta_{nk}x}+D^{\prime}_{nk}e^{-i\eta_{nk}x},\penalty 10000\ x\in\text{D}.$ (11)
We emphasize that in region A, the wavefunction must decay exponentially, as indicated by the Schrödinger equation for a classically forbidden region. This behavior is fundamentally different from the planewave-like character of the wavefunctions in regions B, C, and D. With the eight numerical conditions (six continuity conditions at the points of $x=\{x_{1},x_{2},x_{3}\}$; two periodic boundary conditions at $x=0$ and $x=x_{4}$), the eight coefficients in Eq. (11) can be strictly solved. The potential-barrier affinity effect can occur in region C with planewave-like wavefunctions, as demonstrated in the maintext of our manuscript. We note that systematically adjusting the relative height and spatial separation of the potential barriers in this generalized model may yield deeper insights into the understanding of the formation mechanisms of different bonding types.
Pauli potential
Figure 9: (a) The interacting fermion systems and reference noninteracting systems. The all-electron KSDFT calculations of effective potentials and densities for (b) Na-hP4 along $1\bar{1}$1] direction, (c) FCC-Al along [111] direction, and (d) Diamond along [111] direction.
Although early theoretical models attributed the interstitial-potential well to Pauli exclusion from core electrons[21, 2], no accurate or quantitative theory exists within the fermionic quasiparticle framework to describe the Pauli potential. Here, to comprehensively analyze the impact of Pauli exclusion, we mapped the noninteracting fermions (Kohn-Sham system) to the reference noninteracting bosons, where the bosons have the same charge, mass, point-like shape, chemical potential, and density as the electrons[39, 40], as shown in Fig. 9(a). In KSDFT, the electronic total energy of the interacting systems is defined as:
$E[\rho]=T_{s}[\rho]+E_{H}[\rho]+E_{XC}[\rho]+E_{NE}[\rho],$ (12)
where $E_{H}$, $E_{XC}$, and $E_{NE}$ are the Hartree energy, exchange-correlation energy, and nuclei-electron interacting energy, respectively. With the reference system of noninteracting bosons, the noninteracting kinetic energies ($T_{s}$) can be rewritten as[40]:
$T_{s}[\rho]=T_{vW}[\rho]+T_{P}[\rho],$ (13)
where $T_{vW}[\rho]=-\frac{1}{2}\int{\sqrt{\rho({\mathbf{r}})}\nabla^{2}\sqrt{\rho({\mathbf{r}})}}d^{3}{\mathbf{r}}$ is the von Weizsäcker (vW) kinetic energy[41] for the noninteracting bosons occupying the single collective orbital $\phi_{B}({\mathbf{r}})=\sqrt{\rho({\mathbf{r}})}$ (or Madelung wavefunction[42]). The Pauli kinetic energy ($T_{P}$) is defined as the difference between the vW kinetic energy and the exact one due to the Pauli exclusion. Then, we can derive the effective potential by performing the variation on the total energy functional to be stationary with the conservation of electron number:
$\mu=\frac{\delta E[\rho]}{\delta\rho({\mathbf{r}})}=$ $\frac{\delta T_{v}W[\rho]}{\delta\rho({\mathbf{r}})}+\frac{\delta T_{P}[\rho]}{\delta\rho({\mathbf{r}})}+\frac{\delta(E_{H}[\rho]+E_{XC}[\rho]+E_{NE}[\rho])}{\delta\rho({\mathbf{r}})}$
$=$ $V_{vW}[\rho]({\mathbf{r}})+V_{P}[\rho]({\mathbf{r}})+V_{eff}^{KS}[\rho]({\mathbf{r}}),$ (14)
where $V_{vW}[\rho]({\mathbf{r}})=-\frac{1}{2}{\nabla^{2}\sqrt{\rho({\mathbf{r}})}}/{\sqrt{\rho({\mathbf{r}})}}$ and $V_{eff}^{KS}[\rho]({\mathbf{r}})$ are the vW kinetic potential and Kohn-Sham effective potential. Although the exact and analytical Pauli potential ($V_{P}$) is still unknown, we can numerically calculate the effective potential for the reference bosons using the ground-state electron density and chemical potential calculated from KSDFT:
$V_{eff}^{B}[\rho]({\mathbf{r}})=V_{P}[\rho]({\mathbf{r}})+V_{eff}^{KS}[\rho]({\mathbf{r}})=\mu-V_{vW}[\rho]({\mathbf{r}}),$ (15)
where the chemical potential is set as the Fermi energy level of KSDFT in this work ($\mu=E_{F}$). To include the interaction between nuclear and valence electrons, we performed the all-electron (AE) KSDFT calculations using the full-potential linearised augmented-planewave method[43] by the ELK-v10.6.2 code[44]. The AE effective potentials and electron densities are presented in Figure 9(c-d) for Na-hP4 $\bar{1}1$], Al-FCC [111], and Diamond [111] directions. In this framework, the collective quantum state exhibits bound-state behavior under the all-electron effective potential containing Pauli interaction. The results completely overwhelm the behavior of fermionic quasiparticles, making it exceptionally difficult to analyze electron behavior. Therefore, the PBA effect derived from our approach offers a mathematically rigorous alternative that captures the essential physics. For the AE KSDFT calculations, we applied $18\times 18\times 11$ and $18\times 18\times 18$ $k$-grid meshes for the ground-state calculations of Na-hP4 and Al-FCC/Diamond, respectively. The potential and electron density are represented in the dense real-space grids of $120\times 120\times 120$. The eight-order finite-difference method[45] is adopted for evaluating vW kinetic potential.
Acknowledgements
This research was supported by the National Natural Science Foundation of China under Grants Nos. 12305002, 12304021; the National Key Research and Development Program of China (Grant No. 2023YFA1406200), the Program for JLU Science and Technology Innovative Research Team. Part of the calculation was performed in the high-performance computing center of Jilin University and Ningbo University.
Author contributions statement
Q.X., Z.L., and Y.M. conceived the project. Q.X. designed the theoretical framework and implemented the Kronig-Penney model solver. Q.X. and Z.L. performed the computer simulations. Q.X., Z.L., and Y.M. analyzed the results and wrote the manuscripts. All authors contributed to the discussion and revision of the manuscript.
Additional information
Competing interests The authors declare no competing interests.